OpenStudy (anonymous):

[6.02] For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation?3x + y = 8 -x - 2y = -10

hero (hero):

Hint: Isolate x in the second equation

OpenStudy (anonymous):

so -2y-10

hero (hero):

Sorry, but you didn't isolate x properly. Try again.

OpenStudy (anonymous):

? is that right

hero (hero):

-x - 2y = -10 -(x+2y) = -10 x + 2y = 10 x = 10 - 2y